# Agentic Very Long Video Understanding

### v1

@@ 1 Introduction
Unlocking always-on personal AI assistants requires understanding not just isolated events, but a continuous stream of evolving user experiences. The recent emergence of AI-equipped wearable consumer devices such as the Ray-Ban Meta glasses, Amazon Echo Frames and Snapchat Spectacles as well as various prototypes (engel2023project; xu2025designing) creates an opportunity for AI agents to maintain persistent access to what users see and do over time. For such assistants to provide helpful, personalized, and context-aware assistance, they need to possess a longitudinal video understanding, i.e. recall and interpret a user’s lived experience over extremely long periods of time (days and months).
@@
In this work, we address the challenge of “very long video understanding”. In prior literature, the definition of “long” has been continuously evolving. Popular benchmarks like MSR-VTT (xu2016msrvtt) and DiDeMo (hendricks2017didemo) where videos are up to a minute in length were once considered long, but recent works have further pushed this frontier to several minutes (Wu2024LongVideoBenchAB; grauman2022ego4d) and up to an hour (fu2025videomme; Zhou2024MLVUBM; wang2025lvbench). The recent EgoLife (yang2025egolife) pushes this frontier to 50 hours of Egocentric video over the course of a week, which is the length we define as very long. Unlike previous benchmarks that focus on large numbers of short, independent videos, EgoLife offers continuous, longitudinal first-person video from six individuals. This week-long horizon enables new research directions, such as tracking entities and their interactions across multiple days, analyzing repeated behaviors and habits, and handling extended periods of inactivity or “lulls” in the video stream. Agentic approaches, which equip agents with tools to search, retrieve, and reason over large corpora, have shown potential in addressing some of these limitations (fan2024videoagent; wang2024videoagent; ma2025drvideo; chu2025graphvideoagent). Existing agentic approaches often struggle to maintain coherent reasoning about entities and their relationships over extended temporal horizons, and have difficulty with fine-grained temporal localization such as tracking repeated actions or habits across days (e.g. “how often did I drink water this week?”). Importantly, there is a need for effective linkage between information from different modalities to support richer and more accurate reasoning.
@@
To address these challenges, we propose EGAgent, an enhanced agentic approach that centers on the extraction and use of an entity scene graph from long videos, where nodes represent people, places, and objects, and edges capture their relationships (e.g. uses, interacts with, mentions, talks to). Each node is annotated with temporal intervals indicating when the relation holds. In our system, we equip a planning agent with the ability to search and reason over this entity graph, as well as utilize a visual search tool (SQL + semantic search hybrid) and an audio transcript search tool. As illustrated in figure˜1, the system uses this graph in combination with audio and visual search to locate all shopping-related taxi rides across multiple days and infer who consistently sits next to the user. By leveraging structured representations like entity graphs, our system preserves complex relationships and supports detailed, compositional reasoning over extended timeframes, overcoming the limitations of existing methods.
@@
We evaluate our EGAgent pipeline on the EgoLifeQA benchmark and demonstrate state-of-the-art performance. Notably, EGAgent surpasses the previous state-of-the-art by 32% and 39.7% on the RelationMap and TaskMaster categories respectively, both of which require multi-hop relational reasoning. Our method also achieves competitive results on the Video-MME (Long) benchmark.
@@ 2 Related Work
Video Understanding with Graph-based RAG. Retrieval-augmented generation (RAG) mitigates the context limitations of LLMs by retrieving relevant information from external sources (lewis2020rag; gao2023ragsurvey), which has also been extended to multimodal documents and long-video understanding (yuvis2025rag; faysse2025colpali). Traditional RAG operates over isolated text chunks, often losing relational context. To address this, Graph-based RAG methods such as GraphRAG (edge2024local) and LightRAG (guo2024lightrag) leverages knowledge graphs built from extracted entities and relations from the text corpus. More recently, researchers have begun to explore multi-modal RAG approaches, such as retrieving image frames directly instead of retrieving pre-generated video captions (reddy2025video; wan2025clamr). This approach preserves visual details that may be lost in textual abstraction, enabling more precise and comprehensive responses to complex queries. For instance, Video-RAG (luo2025videorag) performs multi-modal RAG on video frames, automatic speech recognition (ASR) results, optical character recognition (OCR) results, and object-detection results. However, directly retrieving frames also introduces new challenges, including the need for efficient and accurate indexing, retrieval mechanisms, and effective data representations (reddy2025video; wan2025clamr). VideoRAG (ren2025videorag) combines text-, visual-, and graph-based clip retrieval, matching queries to entity descriptions within a graph. AdaVideoRAG (xue2025adavideorag) adaptively selects between no retrieval, naive retrieval, and graph-based retrieval based on question difficulty. RAVU (malik2025ravu) uses VLMs to detect entities, generate frame descriptions, build spatio-temporal graphs, and infer answers. GraphVideoAgent (chu2025graphvideoagent) iteratively retrieves relevant frames via caption-derived graphs. VideoMindPalace (huang2025building) constructs layered spatio-temporal graphs encoding indoor layouts and activity zones, though its reliance on room-level structure limits robustness in open-ended scenes.
@@
Agentic Video Understanding. Recent advances in agentic video understanding have focused on developing systems that can autonomously perceive, reason, and act based on video content (chen2025lvagent). VideoAgent (wang2024videoagent) introduces an agent-based framework where the agent is tasked with iteratively finding the relevant frames in the video for VQA if the information in the initial frames is not sufficient to answer the question. VideoAgent (fan2024videoagent) iteratively employs tools such as object memory search and video-segment search based on video captions and visual embeddings to reach an answer. DrVideo (ma2025drvideo) reframes long-video understanding as long-document understanding by converting videos into text documents, iteratively augmenting them with key frame information and agent-based searches until enough information is gathered for chain-of-thought prediction.
@@ 3 Method
Here we formalize the task of very long video understanding (section˜3.1) and extracting entity graph representations of such long videos (section˜3.2). Lastly, we discuss the design of our proposed agentic framework EGAgent which utilizes these entity graph representations for very long video understanding (section˜3.3).
@@ 3.1 Task Setup
Naively feeding all frames and transcripts into a multimodal LLM or VLM for such very long videos is infeasible due to context window limitations. The prevailing approach, Video Retrieval Augmented Generation (RAG) (luo2025videorag), first selectively retrieves a small subset of frames and audio transcripts deemed relevant to the user query and conditions the VLM on this retrieved set to generate the answer . However, a naive RAG approach over very long egocentric videos is insufficient to answer egocentric queries which are often entity-centric and require multi-hop reasoning across days. These include tracking repeated behaviors, or interactions between specific people, objects, and locations. Direct embedding based retrieval over unstructured clips or captions struggle to maintain coherent entity identities over time to support compositional constraints such as “all times I talked to person X this week”.
@@ 3.2 Entity Graph Representations
To address this, we construct an entity graph to capture relationships and interactions, enabling our planning agent to query this graph during inference.
@@
Edges (): relationships (i.e. interacts with, mentions, talks to, uses), and temporal information
@@ 3.3 Agentic Framework EGAgent
Visual Search Tool samples video frames at 1FPS and embeds each frame as using a vision-encoder (tschannen2025siglip). The generated embeddings along with attributes such as timestamp, location are stored in a vector database which supports efficient retrieval. At inference, the Planning Agent provides a text sub-query (embedded as ) and optional attribute filters (e.g. “kitchen", “morning”). The tool computes cosine similarity for filtered rows in the vector database returning the -nearest neighbors for further analysis.
@@ 4 Experiments
We evaluate EGAgent against baselines on two benchmarks, EgoLifeQA and Video-MME (Long), which focus on very long video understanding. Here we discuss implementation details of our EGAgent and analyze its performance on these datasets. Lastly, we discuss a few ablation studies on entity graph extraction and wall-clock latency.
@@ 4.1 Evaluation Benchmarks
EgoLifeQA: EgoLifeQA consists of 500 long-context Multiple-Choice Questions (MCQs) derived from the EgoLife (yang2025egolife) dataset, in which six participants lived together for one week, continuously recording their daily activities using Project Aria glasses (engel2023project). The benchmark focuses on the 50 hours of videos taken from the perspective of Jake, one of the six participants. The MCQs cover practical questions such as locating items, recalling past events, tracking habits, and analyzing social interactions. Each question has four candidate answers with a single correct option. Each question is associated with query time (e.g. 11:34 AM on day 4) and a manually verified target time, indicating the specific portion of the video that contains the information needed to answer the MCQ correctly.
@@ 4.3 Analysis on EgoLifeQA Benchmark
We compare our approach against various strong baselines in three categories: 1) MLLM with uniform sampling; 2) MLLM with RAG; and 3) existing agentic approaches.
@@
We compare our approach with the following existing agentic methods: EgoButler (yang2025egolife), a hierarchical text-based Retrieval-Augmented Generation (RAG) approach, and Ego-R1 (tian2025egor1), a lightweight 3B-parameter agent trained on egocentric data, including portions of EgoLife for tool calling. We report results of all RAG and existing agentic approaches in table˜1 directly from these works.
Performance Analysis. table˜1 presents a comprehensive comparison of methods on the EgoLifeQA benchmark. Our proposed agentic system, which incorporates entity graph reasoning, achieves strong performance across all evaluation categories and establishes a new state-of-the-art. Notably, while Gemini 2.5 Pro with uniform sampling already outperforms the previous best results (EgoButler), our agentic system based on Gemini 2.5 Pro delivers an additional improvement of 10.7%, highlighting the significant value of entity graph reasoning.
@@
To compare with existing agentic systems, we run our agentic system on the same LLM backbone (GPT-4o) with other agentic system. Our entity graph agent consistently surpasses other agentic approaches utilizing the same model, including EgoButler (+8.4%), VideoAgent (+15.4%), and Ego-R1 (+8.6%).
@@ 4.5 Ablations
Agent Wall-Clock Latency. We tabulate the wall-clock latency of our EGAgent pipeline in table˜4. EGAgent takes between two and three minutes to answer an MCQ, depending on the number of sub-tasks required by the planning agent. We also evaluate the latency impact of the transcript search and replace our default LLM search with BM25 (robertson2009bm25), which drops token usage by at the cost of a 6.8% MCQ accuracy drop on average.

### v2

@@ 1 Introduction
Unlocking always-on personal AI assistants requires understanding not just isolated events, but a continuous stream of evolving user experiences. The recent emergence of AI-equipped wearable consumer devices such as the Ray-Ban Meta glasses, Amazon Echo Frames and Snapchat Spectacles as well as various prototypes (engel2023project; xu2025designing) creates an opportunity for AI agents to maintain persistent access to what users see and do over time. For such assistants to provide helpful, personalized, and context-aware assistance, they need to possess longitudinal video understanding, i.e., the ability to recall and interpret a user’s lived experience over extremely long periods of time (days and months).
@@
In this work, we address the challenge of “very long video understanding”. In prior literature, the definition of “long” has been continuously evolving. Popular benchmarks like MSR-VTT (xu2016msrvtt) and DiDeMo (hendricks2017didemo) where videos are up to a minute in length were once considered long, but recent works have further pushed this frontier to several minutes (Wu2024LongVideoBenchAB; grauman2022ego4d) and up to an hour (fu2025videomme; Zhou2024MLVUBM; wang2025lvbench). The recent EgoLife (yang2025egolife) pushes this frontier to beyond 50 hours of Egocentric video over the course of a week, which is the length we define as very long. Unlike previous benchmarks that focus on large numbers of short, independent videos, EgoLife offers continuous, longitudinal first-person video from six individuals. This week-long horizon enables new research directions, such as tracking entities and their interactions across multiple days, analyzing repeated behaviors and habits, and handling extended periods of inactivity or “lulls” in the video stream. Agentic approaches, which equip agents with tools to search, retrieve, and reason over large corpora, have shown potential in addressing some of these limitations (fan2024videoagent; wang2024videoagent; ma2025drvideo; chu2025graphvideoagent). Existing agentic approaches often struggle to maintain coherent reasoning about entities and their relationships over extended temporal horizons, and have difficulty with fine-grained temporal localization such as tracking repeated actions or habits across days (e.g., “how often did I drink water this week?”). Importantly, effective linkage between information from different modalities is needed to support richer and more accurate reasoning.
@@
To address these challenges, we propose EGAgent, an enhanced agentic approach that centers on the extraction and use of an entity scene graph from long videos, where nodes represent people, places, and objects, and edges capture their relationships (e.g., uses, interacts with, mentions, talks to). Each edge is annotated with temporal intervals indicating when the relation holds. In our proposed EGAgent system, we equip a planning agent with the ability to search and reason over this entity graph, as well as utilize a visual search tool (SQL + semantic search hybrid) and an audio transcript search tool. As illustrated in figure˜1, the system uses this graph in combination with audio and visual search to locate all shopping-related taxi rides across multiple days and infer who consistently sits next to the user. By leveraging structured representations like entity graphs, EGAgent preserves complex relationships and supports detailed, compositional reasoning over extended timeframes, overcoming the limitations of existing methods.
@@
We evaluate our EGAgent pipeline on the EgoLifeQA benchmark and demonstrate state-of-the-art performance. Notably, EGAgent surpasses the previous state-of-the-art by 32% and 39.7% on the RelationMap and TaskMaster categories respectively, both of which require multi-hop relational reasoning. EGAgent also achieves competitive results on the Video-MME (Long) benchmark.
@@ 2 Related Work
Video Understanding with Graph-based RAG. Retrieval-augmented generation (RAG) mitigates the context limitations of LLMs by retrieving relevant information from external sources (lewis2020rag; gao2023ragsurvey), which has also been extended to multimodal documents and long-video understanding (yuvis2025rag; faysse2025colpali). Traditional RAG operates over isolated text chunks, often losing relational context. To address this, Graph-based RAG methods such as GraphRAG (edge2024local) and LightRAG (guo2024lightrag) leverage knowledge graphs built from extracted entities and relations from the text corpus. More recently, researchers have begun to explore multi-modal RAG approaches, such as retrieving image frames directly instead of retrieving pre-generated video captions (reddy2025video; wan2025clamr). This approach preserves visual details that may be lost in textual abstraction, enabling more precise and comprehensive responses to complex queries. For instance, Video-RAG (luo2025videorag) performs multi-modal RAG on video frames, automatic speech recognition (ASR) results, optical character recognition (OCR) results, and object-detection results. However, directly retrieving frames also introduces new challenges, including the need for efficient and accurate indexing, retrieval mechanisms, and effective data representations (reddy2025video; wan2025clamr). VideoRAG (ren2025videorag) combines text-, visual-, and graph-based clip retrieval, matching queries to entity descriptions within a graph. AdaVideoRAG (xue2025adavideorag) adaptively selects between no retrieval, naive retrieval, and graph-based retrieval based on question difficulty. RAVU (malik2026ravu) uses VLMs to detect entities, generate frame descriptions, build spatio-temporal graphs, and infer answers. GraphVideoAgent (chu2025graphvideoagent) iteratively retrieves relevant frames via caption-derived graphs. VideoMindPalace (huang2025building) constructs layered spatio-temporal graphs encoding indoor layouts and activity zones, though its reliance on room-level structure limits robustness in open-ended scenes.
@@
Agentic Video Understanding. Recent advances in agentic video understanding have focused on developing systems that can autonomously perceive, reason, and act based on video content (chen2025lvagent). VideoAgent (wang2024videoagent) introduces an agent-based framework where the agent is tasked with iteratively finding the relevant frames in the video for VQA if the information in the initial frames is not sufficient to answer the question. VideoAgent (fan2024videoagent) iteratively employs tools such as object memory search and video-segment search based on video captions and visual embeddings to reach an answer. DrVideo (ma2025drvideo) reframes long-video understanding as long-document understanding by converting videos into text documents, iteratively augmenting them with key frame information and agent-based searches until enough information is gathered for chain-of-thought prediction. Similarly SiLVR (zhang2026silvr) operates in the text domain by compressing dense visual captions and using a downstream reasoning LLM for video understanding.
@@ 3 Method
Here we formalize the task of very long video understanding (section˜3.1) and extracting entity graph representations of such long videos (section˜3.2). Lastly, we discuss the design of the proposed agentic framework EGAgent which utilizes these entity graph representations for very long video understanding (section˜3.3).
@@ 3.1 Task Setup
Naively feeding all frames and transcripts into a multimodal LLM or VLM for such very long videos is infeasible due to context window limitations. The prevailing approach, Video Retrieval Augmented Generation (RAG) (luo2025videorag), first selectively retrieves a small subset of frames and audio transcripts deemed relevant to the user query and conditions the VLM on this retrieved set to generate the answer . However, a naive RAG approach over very long egocentric videos is insufficient to answer egocentric queries which are often entity-centric and require multi-hop reasoning across days. These include tracking repeated behaviors, or interactions between specific people, objects, and locations. Direct embedding-based retrieval over unstructured clips or captions struggles to maintain coherent entity identities over time to support compositional constraints such as “all times I talked to person X this week”.
@@ 3.2 Entity Graph Representations
To address this, we construct an entity graph to capture relationships and interactions, enabling the planning agent to query this graph during inference.
@@
Edges (): relationships (i.e., interacts with, mentions, talks to, uses), and temporal information
@@ 3.3 Agentic Framework EGAgent
Visual Search Tool samples video frames at 1FPS and embeds each frame as using a vision-encoder (tschannen2025siglip). The generated embeddings along with attributes such as timestamp, location are stored in a vector database which supports efficient retrieval. At inference, the Planning Agent provides a text sub-query (embedded as ) and optional attribute filters (e.g., “kitchen", “morning”). The tool computes cosine similarity for filtered rows in the vector database returning the -nearest neighbors for further analysis.
@@ 4 Experiments
We evaluate EGAgent against baselines on two benchmarks, EgoLifeQA and Video-MME (Long), which focus on very long video understanding. Here we discuss implementation details of EGAgent and analyze its performance on these datasets. Lastly, we discuss a few ablation studies on entity graph extraction and wall-clock latency.
@@ 4.1 Evaluation Benchmarks
EgoLifeQA: EgoLifeQA consists of 500 long-context Multiple-Choice Questions (MCQs) derived from the EgoLife (yang2025egolife) dataset, in which six participants lived together for one week, continuously recording their daily activities using Project Aria glasses (engel2023project). The benchmark focuses on the 50 hours of videos taken from the perspective of Jake, one of the six participants. The MCQs cover practical questions such as locating items, recalling past events, tracking habits, and analyzing social interactions. Each question has four candidate answers with a single correct option. Each question is associated with query time (e.g., 11:34 AM on day 4) and a manually verified target time, indicating the specific portion of the video that contains the information needed to answer the MCQ correctly.
@@ 4.3 Analysis on EgoLifeQA Benchmark
We compare the proposed EGAgent against strong baselines in three categories: 1) MLLM with uniform sampling; 2) MLLM with RAG; and 3) existing agentic approaches.
@@
We compare our EGAgent with the following existing agentic methods: EgoButler (yang2025egolife), a hierarchical text-based Retrieval-Augmented Generation (RAG) approach, and Ego-R1 (tian2025egor1), a lightweight 3B-parameter agent trained on egocentric data, including portions of EgoLife for tool calling. We report results of all RAG and existing agentic approaches in table˜1 directly from these works.
Performance Analysis. table˜1 presents a comprehensive comparison of methods on the EgoLifeQA benchmark. Our proposed agentic system, which incorporates entity graph reasoning, achieves strong performance across all evaluation categories and establishes a new state-of-the-art. Notably, while Gemini 2.5 Pro with uniform sampling already outperforms the previous best results (EgoButler), our EGAgent with a Gemini 2.5 Pro backbone delivers an additional improvement of 10.7%, highlighting the significant value of entity graph reasoning.
@@
To compare with existing agentic systems, we run EGAgent on the same LLM backbone (GPT-4o) with other agentic system. Our entity graph agent consistently surpasses other agentic approaches utilizing the same model, including EgoButler (+8.4%), VideoAgent (+15.4%), and Ego-R1 (+8.6%).
@@ 4.5 Ablations
Agent Wall-Clock Latency. We tabulate the wall-clock latency of EGAgent pipeline in table˜4. EGAgent takes between two and three minutes to answer an MCQ, depending on the number of sub-tasks required by the planning agent. We also evaluate the latency impact of the transcript search and replace the default LLM search with BM25 (robertson2009bm25), which drops token usage by at the cost of a 6.8% MCQ accuracy drop on average.